OpenStudy (anonymous):

how to find the LCD of 12, 9, 60, and 45?

OpenStudy (mathhelp346):

multiply them all

OpenStudy (mathhelp346):

12*9*60*45

OpenStudy (anonymous):

o-o okkaaayyy i dont think thats how u do it, tho.....

OpenStudy (mathhelp346):

is 12*9*60*5 possible?

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

but that is not how u do it....

OpenStudy (mathhelp346):

pfft

OpenStudy (anonymous):

XD srry

OpenStudy (mathhelp346):

lol

jimthompson5910 (jim_thompson5910):

12 = 2*2*3 9 = 3*3 60 = 2*2*3*5 45 = 3*3*5 Now highlight all unique terms that occur the most. These terms are 2*2, 3*3 and 5 Multiply them together to get: 2*2*3*3*5 = 4*9*5 = 36*5 = 180 So the LCM of those numbers is 180